OpenStudy (anonymous):

given f(x)=sqrt{6x} and g(x)=1/x-6 . which value is in the domain 8 4 2 6 ?

OpenStudy (fifciol):

\[6x \ge 0\] and \[x-6\neq 0\] so \[x \ge 0\] and \[x \neq 6\] so \[D: x \in <0;6) \cup (6: \infty)\]

OpenStudy (fifciol):

is that what you mean?

OpenStudy (anonymous):

it was wrong *tear*